If a plane can travel 470 miles per hour with the wind and 410 miles per hour against the wind, find the speed of the plane in the still air?
Answer by mananth(16949)

Ilet x be the speed of the plane in still air
y be the speed of the wind
Speed of plane with wind = x + y
Speed of plane against wind= x - y.
So we have the following system of equations:
{ x + y = 470. . . . .1
{ x - y = 410. . . . 2
Adding both we get 2x = 880
/2
x = 440.
Therefore, the speed of the plane in still air is 440mph.
...
m.ananth@hotmail.ca